# MSDNet: Multi-Scale Decoder for Few-Shot Semantic Segmentation via Transformer-Guided Prototyping

### v1

@@ III-B4 Multi Scale Decoder
III-B5 Spatial Transformer Decoder (STD)
@@
III-C Loss function
@@
TABLE I: Performance on $PASCAL-5^{i}$ in terms of mIoU and FB-IoU. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@
PANet [39] ICCV19 44.0 57.5 50.8 44.0 49.1 - 55.3 67.2 61.3 53.2 59.3 - 23.5M
PGNet [38] ICCV19 56.0 66.9 50.6 50.4 56.0 69.9 57.7 68.7 52.9 54.6 58.5 70.5 17.2M
PFENet [47] TPAMI20 61.7 69.5 55.4 56.3 60.8 73.3 63.1 70.7 55.8 57.9 61.9 73.9 10.3M
PPNet [48] ECCV20 48.6 60.6 55.7 46.5 52.8 69.2 58.9 68.3 66.8 58.0 63.0 75.8 31.5M
RePRI [40] CVPR21 59.8 68.3 62.1 48.5 59.7 - 64.6 71.4 71.1 59.3 66.6 - -
HSNet [43] ICCV21 64.3 70.7 60.3 60.5 64 76.7 70.3 73.2 67.4 67.1 69.5 80.6 2.5M
CyCTR [44] NeurIPS21 65.7 71.0 59.5 59.7 64.0 - 69.3 73.5 63.8 63.5 67.5 - 15.4M
NTRENet [49] CVPR22 65.4 72.3 59.4 59.8 64.2 77.0 66.2 72.8 61.7 62.2 65.7 78.4 19.9M
ABCNet [50] CVPR23 62.5 70.8 57.2 58.1 62.2 74.1 64.7 73.0 57.1 59.5 63.6 74.2 -
QGPLNet [51] ACM TOMM23 56.95 68.99 60.1 54.98 60.25 - 61.78 70.96 69.56 58.26 65.14 - -
DRNet [52]
@@
FWB [53] ICCV19 51.3 64.5 56.7 52.2 56.2 - 54.8 67.4 62.2 55.3 59.9 - 43.0M
PPNet [48] ECCV20 52.7 62.8 57.4 47.7 55.2 70.9 60.3 70.0 69.4 60.7 65.1 77.5 50.5M
PFENet [47] TPAMI20 60.5 69.4 54.4 55.9 60.1 72.9 62.8 70.4 54.9 57.6 61.4 73.5 10.3M
HSNet [43] ICCV21 67.3 72.3 62.0 63.1 66.2 77.6 71.8 74.4 67.0 68.3 70.4 80.6 2.5M
CWT [41] ICCV21 56.9 65.2 61.2 48.8 58 - 62.6 70.2 68.8 57.2 64.7 - -
CyCTR [44] NeurIPS21 69.3 72.7 56.5 58.6 64.3 73.0 73.5 74.0 58.6 60.2 66.6 75.4 15.4M
NTRENet [49] CVPR22 65.5 71.8 59.1 58.3 63.7 75.3 67.9 73.2 60.1 66.8 67.0 78.2 19.9M
ABCNet [50] CVPR23 62.7 70.0 55.1 57.5 61.3 73.7 63.4 71.8 56.4 57.7 62.3 74 -
QGPLNet [51] ACM TOMM23 59.66 69.77 65.15 55.9 62.64 - 65.05 72.75 71.12 59.85 67.19 - -
DRNet [52]
@@
TABLE II: Performance on $COCO-20^{i}$ in terms of mIoU and FB-IoU. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@
ResNet50 PPNet [48] ECCV20 28.1 30.8 29.5 27.7 29.0 - 39.0 40.8 37.1 37.3 38.5 - 31.5M
PFENet [47] TPAMI20 36.5 38.6 34.5 33.8 35.8 - 36.5 43.3 37.8 38.4 39.0 - 10.3M
HSNet [43] ICCV21 36.3 43.1 38.7 38.7 39.2 68.2 43.3 51.3 48.2 45.0 46.9 70.7 2.5M
CyCTR [44] NeurIPS21 38.9 43.0 39.6 39.8 40.3 - 41.1 48.9 45.2 47.0 45.6 - 15.4M
NTRENet [49] CVPR22 36.8 42.6 39.9 37.9 39.3 68.5 38.2 44.1 40.4 38.4 40.3 69.2 19.9M
BAM [54] CVPR22 43.4 50.6 47.5 43.4 46.2 - 49.3 54.2 51.6 49.6 51.2 - 26.7M
DCAMA [15] ECCV22 41.9 45.1 44.4 41.7 43.3 69.5 45.9 50.5 50.7 46.0 48.3 71.7 47.7M
ABCNet [50] CVPR23 36.5 35.7 34.7 31.4 34.6 59.2 40.1 40.1 39.0 35.9 38.8 62.8 -
DRNet [52]
@@
PFENet++ [56] TPAMI24 40.9 46.0 42.3 40.1 42.3 65.7 47.5 53.3 47.3 46.4 48.6 70.3 -
@@
ResNet101 FWB [53] ICCV19 17.0 18.0 21.0 28.9 21.2 - 19.1 21.5 23.9 30.1 23.7 - 43.0M
PFENet [47] TPAMI20 36.8 41.8 38.7 36.7 38.5 63.0 40.4 46.8 43.2 40.5 42.7 65.8 10.3M
HSNet [43] ICCV21 37.2 44.1 42.4 41.3 41.2 69.1 45.9 53.0 51.8 47.1 49.5 72.4 2.5M
NTRENet [49] CVPR22 38.3 40.4 39.5 38.1 39.1 67.5 42.3 44.4 44.2 41.7 43.2 69.6 19.9M
DCAMA [15] ECCV22 41.5 46.2 45.2 41.3 43.5 69.9 48.0 58.0 54.3 47.1 51.9 73.3 47.7M
ABCNet [50] CVPR23 40.7 45.9 41.6 40.6 42.2 66.7 43.2 50.8 45.8 47.1 46.7 62.8 -
DRNet [52]
@@
PFENet++ [56] TPAMI24 42.0 44.1 41.0 39.4 41.6 65.4 47.3 55.1 50.1 50.1 50.7 70.9 -
@@
IV Experimental Results
IV-A Datasets
We evaluated our proposed method on two widely used datasets commonly employed in few-shot segmentation tasks: $PASCAL-5^{i}$ [57] and $COCO-20^{i}$ [53].
[…]
IV-B Experimental Setting
IV-C Evaluation Metrics
@@
IV-D Comparison with SOTA
In this subsection, we compare our proposed method with several SOTA methods on both the $PASCAL-5^{i}$ and $COCO-20^{i}$ datasets. We present the results in Table I and Table II, respectively, where we report the mIoU and FB-IoU scores under both 1-shot and 5-shot settings, along with the final FB-IoU value. The results of other methods are obtained from their respective original papers.
@@ IV-D Comparison with SOTA
It is important to highlight that our proposed method maintains a remarkably low number of learnable parameters, with only 1.5 million parameters. This stands in stark contrast to some SOTA methods, which possess significantly higher parameter counts, exceeding 40 million parameters in certain cases. This demonstrates the efficiency and effectiveness of our approach in achieving superior segmentation performance while maintaining a compact model architecture.
IV-E Cross-dataset task
TABLE III: Few-shot segmentation performance on cross-dataset task, ”$COCO-20^{i}\rightarrow PASCAL-5^{i}$”, in terms of mIoU, with different backbones (ResNet-50 and ResNet-101). Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@ IV-E Cross-dataset task
ResNet50 PFENet [47] TPAMI20 43.2 65.1 66.6 69.7 61.1 45.1 66.8 68.5 73.1 63.4
RePRI [40] CVPR21 52.2 64.3 64.8 71.6 63.2 56.5 68.2 70.0 76.2 67.7
HSNet [43] ICCV21 45.4 61.2 63.4 75.9 61.6 56.9 65.9 71.3 80.8 68.7
HSNet-HM [61] ECCV22 43.4 68.2 69.4 79.9 65.2 50.7 71.4 73.4 83.1 69.7
VAT-HM [61] ECCV22 68.3 64.9 67.5 79.8 65.1 55.6 68.1 72.4 82.8 69.7
RTD [62] ECCV22 57.4 62.2 68.0 74.8 65.6 65.7 69.7 70.8 75.0 70.1
PMNet [9] WACV24 68.8 70.0 65.1 62.3 66.6 73.9 74.5 73.3 72.1 73.4
@@
ResNet101 HSNet [43] ICCV21 47.0 65.2 67.1 77.1 64.1 57.2 69.5 72.0 82.4 70.3
HSNetT-HM [61] ECCV22 46.7 68.6 71.1 79.7 66.5 53.7 70.7 75.2 83.9 70.9
RTD [62] ECCV22 59.4 64.3 70.8 72.0 66.6 67.2 72.7 72.0 78.9 72.7
PMNet [9] WACV24 71.0 72.3 66.6 63.8 68.4 75.2 76.3 77.0 72.6 75.3
@@
Figure 5: Qualitative comparison of component effects on $COCO-20^{i}$ dataset in 1-shot scenario

### v3

@@ 3.2.4 Multi Scale Decoder
The multi-scale decoder is primarily responsible for spatially refining the segmentation mask by integrating hierarchical feature information. While earlier modules such as CMGM provide a semantic prior for object localization, the multi-scale decoder enhances boundary precision and structural integrity. By leveraging support features at multiple resolutions and combining them through residual connections, the decoder progressively improves the segmentation quality across scales, enabling more detailed and accurate mask reconstruction.
3.2.5 Spatial Transformer Decoder (STD)
@@
STD serves as a semantic refinement engine that integrates the information distilled by CMGM and complements the spatial reconstruction performed by the multi-scale decoder. By attending to the query features in relation to the contextual support prototypes, STD produces dynamic kernels that capture higher-order dependencies. The resulting semantic-aware embedding is then merged with the output of the multi-scale decoder, allowing the final segmentation prediction to benefit from both semantic precision and spatial detail. This fusion ensures that the strengths of both decoding strategies are harmonized in the final mask generation.
3.3 Loss function
@@
4 Experimental Results
4.1 Datasets
We evaluated our proposed method on two widely used datasets commonly employed in few-shot segmentation tasks: $PASCAL-5^{i}$ shaban2017one and $COCO-20^{i}$ nguyen2019feature .
[…]
4.2 Experimental Setting
Table 1: Performance on $PASCAL-5^{i}$ in terms of mIoU and FB-IoU. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@ 4.2 Experimental Setting
PANet wang2019panet ICCV19 44.0 57.5 50.8 44.0 49.1 - 55.3 67.2 61.3 53.2 59.3 - 23.5M
PGNet zhang2019pyramid ICCV19 56.0 66.9 50.6 50.4 56.0 69.9 57.7 68.7 52.9 54.6 58.5 70.5 17.2M
PFENet tian2020prior TPAMI20 61.7 69.5 55.4 56.3 60.8 73.3 63.1 70.7 55.8 57.9 61.9 73.9 10.3M
PMM yang2020prototype ECCV20 52.0 67.5 51.5 49.8 55.2 - 55.0 68.2 52.9 51.1 56.8 - -
PPNet liu2020part ECCV20 48.6 60.6 55.7 46.5 52.8 69.2 58.9 68.3 66.8 58.0 63.0 75.8 31.5M
RePRI boudiaf2021few CVPR21 59.8 68.3 62.1 48.5 59.7 - 64.6 71.4 71.1 59.3 66.6 - -
ASR liu2021anti CVPR21 55.2 70.3 53.3 53.6 58.1 - 58.3 71.8 56.8 55.7 60.9 - -
SAGNN xie2021scale CVPR21 64.7 69.6 57.0 57.2 62.1 73.2 64.9 70.0 57.0 59.3 62.8 73.3 -
HSNet min2021hypercorrelation ICCV21 64.3 70.7 60.3 60.5 64 76.7 70.3 73.2 67.4 67.1 69.5 80.6 2.5M
CWT lu2021simpler ICCV21 56.3 62.0 59.9 47.2 56.4 - 61.3 68.5 68.5 56.6 63.7 - -
CyCTR zhang2021few NeurIPS21 65.7 71.0 59.5 59.7 64.0 - 69.3 73.5 63.8 63.5 67.5 - 15.4M
NTRENet liu2022learning CVPR22 65.4 72.3 59.4 59.8 64.2 77.0 66.2 72.8 61.7 62.2 65.7 78.4 19.9M
ABCNet wang2023rethinking CVPR23 62.5 70.8 57.2 58.1 62.2 74.1 64.7 73.0 57.1 59.5 63.6 74.2 -
SRPNet ding2023self
Pattern
Recognition23
62.8 69.3 55.8 58.1 61.5 - 64.3 70.3 55.1 60.5 62.6 - -
QGPLNet tang2023query ACM TOMM23 56.95 68.99 60.1 54.98 60.25 - 61.78 70.96 69.56 58.26 65.14 - -
NSF lu2023prediction
IEEE Trans.
img. proc.23
51.8 55.4 50.6 36.9 48.7 - 59.0 64.0 62.7 48.3 58.5 - -
PCN lu2023prediction
IEEE Trans.
img. proc.23
47.9 51.2 51.2 41.3 47.9 - 53.0 58.0 61.6 51.6 56.0 - -
SRPNet ding2023self
Pattern
Recognition23
62.8 69.3 55.8 58.1 61.5 - 64.3 70.3 55.1 60.5 62.6 - -
DRNet chang2024drnet
@@
AFANet ma2025afanet
IEEE Trans.
Multimedia25
65.7 68.7 60.6 61.5 64.0 - 69.0 70.4 61.3 64.0 66.2 - -
MFIRNet chen2025mask Neurocomp.25 65.7 69.2 54.5 49.3 59.7 70.4 - - - - - - -
ESGP zhang2025efficient
Pattern
Recognition25
63.9 72.6 57.1 61.4 63.8 - - - - - - - -
@@
FWB nguyen2019feature ICCV19 51.3 64.5 56.7 52.2 56.2 - 54.8 67.4 62.2 55.3 59.9 - 43.0M
PPNet liu2020part ECCV20 52.7 62.8 57.4 47.7 55.2 70.9 60.3 70.0 69.4 60.7 65.1 77.5 50.5M
DAN wang2020few ECCV20 54.7 68.6 57.8 51.6 58.2 71.9 57.9 69.0 60.1 54.9 60.5 72.3 -
PFENet tian2020prior TPAMI20 60.5 69.4 54.4 55.9 60.1 72.9 62.8 70.4 54.9 57.6 61.4 73.5 10.3M
RePRI boudiaf2021few CVPR21 59.6 68.6 62.2 47.2 59.4 - 66.2 71.4 67.0 57.7 65.6 - -
HSNet min2021hypercorrelation ICCV21 67.3 72.3 62.0 63.1 66.2 77.6 71.8 74.4 67.0 68.3 70.4 80.6 2.5M
CWT lu2021simpler ICCV21 56.9 65.2 61.2 48.8 58 - 62.6 70.2 68.8 57.2 64.7 - -
CyCTR zhang2021few NeurIPS21 69.3 72.7 56.5 58.6 64.3 73.0 73.5 74.0 58.6 60.2 66.6 75.4 15.4M
NTRENet liu2022learning CVPR22 65.5 71.8 59.1 58.3 63.7 75.3 67.9 73.2 60.1 66.8 67.0 78.2 19.9M
ABCNet wang2023rethinking CVPR23 62.7 70.0 55.1 57.5 61.3 73.7 63.4 71.8 56.4 57.7 62.3 74 -
QGPLNet tang2023query ACM TOMM23 59.66 69.77 65.15 55.9 62.64 - 65.05 72.75 71.12 59.85 67.19 - -
NSF lu2023prediction
IEEE Trans.
img. proc.23
52.6 61.9 58.7 41.5 53.7 - 59.9 67.3 65.6 50.4 60.8 - -
DRNet chang2024drnet
@@
Table 2: Performance on $COCO-20^{i}$ in terms of mIoU and FB-IoU. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@
ResNet50 PPNet liu2020part ECCV20 28.1 30.8 29.5 27.7 29.0 - 39.0 40.8 37.1 37.3 38.5 - 31.5M
PMM yang2020prototype ECCV20 29.3 34.8 27.1 27.3 29.6 - 33.0 40.6 30.3 33.3 34.3 - -
RPMM yang2020prototype ECCV20 29.5 36.8 28.9 27.0 30.6 - 33.8 42.0 33.0 33.3 35.5 - -
PFENet tian2020prior TPAMI20 36.5 38.6 34.5 33.8 35.8 - 36.5 43.3 37.8 38.4 39.0 - 10.3M
RePRI boudiaf2021few CVPR21 32.0 38.7 32.7 33.1 34.1 - 39.3 45.4 39.7 41.8 41.6 - -
HSNet min2021hypercorrelation ICCV21 36.3 43.1 38.7 38.7 39.2 68.2 43.3 51.3 48.2 45.0 46.9 70.7 2.5M
CWT lu2021simpler ICCV21 32.2 36.0 31.6 31.6 32.9 - 40.1 43.8 39.0 42.4 41.3 - -
CyCTR zhang2021few NeurIPS21 38.9 43.0 39.6 39.8 40.3 - 41.1 48.9 45.2 47.0 45.6 - 15.4M
NTRENet liu2022learning CVPR22 36.8 42.6 39.9 37.9 39.3 68.5 38.2 44.1 40.4 38.4 40.3 69.2 19.9M
BAM lang2022learning CVPR22 43.4 50.6 47.5 43.4 46.2 - 49.3 54.2 51.6 49.6 51.2 - 26.7M
DCAMA shi2022dense ECCV22 41.9 45.1 44.4 41.7 43.3 69.5 45.9 50.5 50.7 46.0 48.3 71.7 47.7M
ABCNet wang2023rethinking CVPR23 36.5 35.7 34.7 31.4 34.6 59.2 40.1 40.1 39.0 35.9 38.8 62.8 -
DRNet chang2024drnet
@@
PFENet++ luo2023pfenet++ TPAMI24 40.9 46.0 42.3 40.1 42.3 65.7 47.5 53.3 47.3 46.4 48.6 70.3 -
DCP lang2024few
Int. Jour.
Comp. Vision24
43.0 48.6 45.4 44.8 45.5 - 47.0 54.7 51.7 50.0 50.9 - 11.3
PMNet chen2024pixel WACV24 39.8 41.0 40.1 40.7 40.4 - 50.1 51.0 50.4 49.6 50.3 - -
RiFeNet bao2024relevant AAAI24 39.1 47.2 44.6 45.4 44.1 - 44.3 52.4 49.3 48.4 48.6 - -
HSRap luo2025combining
Exp. System
with App.25
43.1 48.5 42.9 41.1 43.8 - 49.2 58.1 52.9 49.9 52.5 - -
AFANet ma2025afanet
IEEE Trans.
Multimedia25
40.2 45.1 44.0 45.1 43.6 - 41.0 49.5 43.0 46.9 45.1 - -
@@
ResNet101 FWB nguyen2019feature ICCV19 17.0 18.0 21.0 28.9 21.2 - 19.1 21.5 23.9 30.1 23.7 - 43.0M
PFENet tian2020prior TPAMI20 36.8 41.8 38.7 36.7 38.5 63.0 40.4 46.8 43.2 40.5 42.7 65.8 10.3M
HSNet min2021hypercorrelation ICCV21 37.2 44.1 42.4 41.3 41.2 69.1 45.9 53.0 51.8 47.1 49.5 72.4 2.5M
CWT lu2021simpler ICCV21 30.3 36.6 30.5 32.2 32.4 - 38.5 46.7 39.4 43.2 42.0 - -
NTRENet liu2022learning CVPR22 38.3 40.4 39.5 38.1 39.1 67.5 42.3 44.4 44.2 41.7 43.2 69.6 19.9M
DCAMA shi2022dense ECCV22 41.5 46.2 45.2 41.3 43.5 69.9 48.0 58.0 54.3 47.1 51.9 73.3 47.7M
ABCNet wang2023rethinking CVPR23 40.7 45.9 41.6 40.6 42.2 66.7 43.2 50.8 45.8 47.1 46.7 62.8 -
QGPLNet tang2023query ACM TOMM23 34.86 40.14 35.68 36.32 36.75 - 42.69 48.94 42.98 43.69 44.58 - -
DRNet chang2024drnet
@@
PFENet++ luo2023pfenet++ TPAMI24 42.0 44.1 41.0 39.4 41.6 65.4 47.3 55.1 50.1 50.1 50.7 70.9 -
PMNet chen2024pixel WACV24 44.7 44.3 44.0 41.8 43.7 - 52.6 53.3 53.5 52.8 53.1 - -
HSRap luo2025combining
Exp. System
with App.25
42.0 50.0 43.5 43.8 44.8 - 50.3 60.1 53.4 50.9 53.9 - -
@@
4.3 Evaluation Metrics
@@
4.4 Comparison with SOTA
In this subsection, we compare our proposed method with several SOTA methods on both the $PASCAL-5^{i}$ and $COCO-20^{i}$ datasets. We present the results in Table 1 and Table 2, respectively, where we report the mIoU and FB-IoU scores under both 1-shot and 5-shot settings, along with the final FB-IoU value. The results of other methods are obtained from their respective original papers.
@@ 4.4 Comparison with SOTA
Our proposed MSDNet consistently performs well across diverse folds and datasets. We observe particularly strong gains in folds with higher object complexity and background clutter (e.g., fold0 and fold1 in both datasets), indicating the effectiveness of the proposed method in capturing fine-grained semantic relations between support and query images.
Compared to heavier models such as HSNet min2021hypercorrelation and DRNet chang2024drnet , MSDNet maintains competitive or superior performance while using significantly fewer parameters. This efficiency stems from our lightweight Transformer-guided decoding strategy and the integration of multi-scale features, which compensates for reduced model size.
In scenarios involving complex object shapes or fine structures, our multi-scale decoder helps refine the mask resolution, especially under the 5-shot setting. However, in some folds with low inter-class variability or where object localization is less ambiguous, larger models with more attention heads (e.g., DCAMA shi2022dense ) may achieve slightly better results due to their higher modeling capacity.
These findings highlight that MSDNet is especially effective in few-shot settings where efficiency, generalization, and contextual matching are critical, offering a strong balance between accuracy and computational cost.
Our method is designed with computational efficiency in mind. MSDNet contains only 1.5 million learnable parameters, which is significantly fewer than many recent few-shot segmentation models. This lightweight design is particularly beneficial for deployment in real-world scenarios where memory and computational resources are limited.
Although MSDNet integrates two decoding branches—namely the Multi-Scale Decoder and the STD—the architectural design remains computationally tractable. The multi-scale decoder is composed of shallow residual blocks and convolutional upsampling, while the STD is implemented using single cross-attention block rather than a deep transformer stack. This careful design ensures that complexity does not grow excessively, even as the model benefits from richer multi-scale and semantic context.
Furthermore, the low parameter count is achieved without sacrificing segmentation accuracy, as demonstrated in our experimental results. This balance between performance and model efficiency makes MSDNet well-suited for practical applications in environments with constrained compute budgets.
Table 3: Few-shot segmentation performance on cross-dataset task, ”$COCO-20^{i}\rightarrow PASCAL-5^{i}$”, in terms of mIoU, with different backbones (ResNet-50 and ResNet-101). Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@
ResNet50 PFENet tian2020prior TPAMI20 43.2 65.1 66.6 69.7 61.1 45.1 66.8 68.5 73.1 63.4
RePRI boudiaf2021few CVPR21 52.2 64.3 64.8 71.6 63.2 56.5 68.2 70.0 76.2 67.7
HSNet min2021hypercorrelation ICCV21 45.4 61.2 63.4 75.9 61.6 56.9 65.9 71.3 80.8 68.7
VAT hong2022cost ECCV22 52.1 64.1 67.4 74.2 64.5 58.5 68.0 72.5 79.9 69.7
HSNet-HM liu2022few ECCV22 43.4 68.2 69.4 79.9 65.2 50.7 71.4 73.4 83.1 69.7
VAT-HM liu2022few ECCV22 68.3 64.9 67.5 79.8 65.1 55.6 68.1 72.4 82.8 69.7
RTD wang2022remember ECCV22 57.4 62.2 68.0 74.8 65.6 65.7 69.7 70.8 75.0 70.1
PMNet chen2024pixel WACV24 68.8 70.0 65.1 62.3 66.6 73.9 74.5 73.3 72.1 73.4
@@
ResNet101 HSNet min2021hypercorrelation ICCV21 47.0 65.2 67.1 77.1 64.1 57.2 69.5 72.0 82.4 70.3
HSNetT-HM liu2022few ECCV22 46.7 68.6 71.1 79.7 66.5 53.7 70.7 75.2 83.9 70.9
RTD wang2022remember ECCV22 59.4 64.3 70.8 72.0 66.6 67.2 72.7 72.0 78.9 72.7
PMNet chen2024pixel WACV24 71.0 72.3 66.6 63.8 68.4 75.2 76.3 77.0 72.6 75.3
@@
4.5 Cross-dataset task